A system of 1520 particles, each of which is either an electron or a proton, has a net charge of −6.784×10−17 C .

How many electrons are in this system?

What is the mass of this system?

Thank you in advance
Answer #1

a) the net charge on the system =-6.784*10-17C

since electron is negatively charged we have=-6.784*10-17/-1.6*10-19c=424 more eletrons

so we can see the number of electrons is higher, so the exact number of protons is=1520-424/2=548

so exact number of electrons=1520-548=972

so the answer is 972 electrons

b) the mass of the system

n1*mass of electron+n2*mass of proton

972*9.11*10-31kg+548 *1.67*10-27kg=9.17*10-25kg

so the answer is 9.17*10-25kg or 9.20*10-25kg